# Annamaria Porreca Fabrizio Maturo Viviana Ventre

### v1

@@ 5 Fuzzy Graphs: Generalization of Weighted Graphs
Assume $G_{dn}=(V,E_{dn})$ is a DFSN,
$\widetilde{C}_{d}=\left(\begin{array}[]{ccc}\widetilde{s}\left(v_{1},v_{1}\right)&\cdots&\widetilde{s}\left(v_{1},v_{n}\right)\\ \vdots&\ddots&\vdots\\ \widetilde{s}\left(v_{n},v_{1}\right)&\cdots&\widetilde{s}\left(v_{n},v_{n}\right)\end{array}\right)$
then $\widetilde{C}_{d}$ is called the directed fuzzy connected intensity matrix of $\widetilde{G}_{dn}$.
The directed fuzzy connected intensity matrix is a crucial concept in DFSN. With it, we can determine the directed fuzzy connected intensity between any two actors in DFSN, indicating the relationship (direct or indirect) between them.
Hereafter, we will focus only on triangular fuzzy numbers using a simple notation like the following:
$\tilde{A}=(a_{l},a_{m},a_{r}).$
An example of FG can be done to represent relations in a social network like Facebook or in a University Department, where the nodes are the department members, and the edges are the relations between them assessed according to a questionnaire. If we collect information about the relationship between pairs of Professors and researchers according to how much they feel they have relations with another member (for example, not at all, a little, a lot, etc.), summarizing this information with a scalar, as we do in the WG, would lead to a loss of information. Indeed, when dealing with this type of opinion, mainly when it is turned off through a human language attribute, the data always has a degree of inaccuracy. In such a context, it is natural to consider that the FG is direct because the opinion of member $A$ about his relationship with member $B$ may differ from that of member B about his relationship with member $A$. For this reason, we consider only the case of direct graphs. We will have the symmetrical FG be only a particular case of the asymmetrical one if all couples give the same opinion. This latter case would be improbable since we deal with personal opinions.
Assuming that through a fuzzy inference procedure, we have translated the qualitative answers of the department members into fuzzy numbers, we could have a situation like the one illustrated in Figure 2.
Figure 2: Example of fuzzy social network representing relationship in a University Department. The edges are represented via triangular fuzzy numbers. For simplicity, we do not plot the fuzzy number when there is a null relationship.
6 Fuzzy Centrality Indexes in a Fuzzy Social Network
Centrality indices quantify the importance of a node in a given network, which is often identified with the importance of the corresponding entity in the complex system modelled by the network. In the following sections, we extend the classical centrality indexes to the case of a network where the edges are represented by triangular fuzzy numbers. In other words, we define fuzzy degree centrality, fuzzy betweenness centrality, fuzzy closeness centrality, and fuzzy eigenvector centrality.
@@ 6.1 Fuzzy degree centrality
In real-world interactions, we often consider people with many connections to be important. Degree centrality transfers the same idea into a measure. The degree centrality measure ranks nodes with more connections higher in terms of centrality. In the classical framework, the degree centrality $C_{d}(v_{i})$, in directed graphs can be computed in different ways according to the number of in-connections and out-connections, leading to the in-degree index $d_{i}^{in}$ and the out-degree index $d_{i}^{out}$, respectively, or the combination of both. However, in the context of (directed) fuzzy social networks, the edges are represented by fuzzy numbers, and thus they are not scalars. Therefore, in the next Sections, we extend the concepts of in-degree centrality and out-degree centrality to the case of fuzzy numbers.
The first step is introducing the concepts of fuzzy triangular ordered weighted arithmetic operator [37]. The Fuzzy Triangular Ordered Weighted Arithmetic Operator (FTOWA) can be defined as follows.
@@
The fuzzy triangular ordered weighted arithmetic operator, denoted by FTOWA, is the mapping FTOWA: $\widetilde{\mathscr{P}}([0,1])^{n}\rightarrow\widetilde{\mathscr{P}}([0,1])$, such that
$\operatorname{FTOWA}\left(\tilde{x}_{1},\ldots,\tilde{x}_{n}\right)=\bigoplus_{i\in[n]}\left(w_{i}\odot\tilde{x}_{(i)}\right)$
where $\oplus$ is a t-conorm, $\odot$ is a t-norm, $\left(w_{1},\ldots,w_{n}\right)$ is a weighting vector associated with $\oplus$, and ()($\cdot$)( ⋅ ) is a permutation on $[n]$ such that $\tilde{x}_{(1)}\leq\cdots\leq\tilde{x}_{(n)}$
@@ 6.1.1 Fuzzy in-degree centrality index
$\widetilde{D}_{in}\left(v_{i}\right)=\bigoplus_{j\in[n],j\neq i}\left(w_{j}\odot\tilde{O}_{j}^{in}\right).$ (26)
To rank the fuzzy numbers, we can use one of the approaches presented in Section 2.12. In the following, we will focus on the centre of gravity approach to understand the ranking of fuzzy numbers.
It follows that we can obtain the following extreme cases:
  • 1.
@@
The index just illustrated in Equation 26 does not allow us to compare the centrality values between different networks. To overcome this drawback, we can normalize the index so that it becomes a pure number between 0 and 1. The reasoning behind the normalization is that we calculate the maximum that the in-degree centrality index can take. The maximum of this index would, of course, be obtained if an actor is perfectly connected with all the other nodes. For this aim, the edge represented by the fuzzy number should have the characteristic of possessing a core with the maximum possible value on the support of the fuzzy number (with zero fuzziness). If the latter happens for all the actor’s edges, then we can compute the maximum of Equation 26 as follows.
The easiest way to find this relative index is to normalize the fuzzy numbers. In fact, unlike the classical case in which the number of edges is necessary to normalize the in-degree centrality, in this case, all the nodes are connected and therefore, what matters is the intensity of the edges. Once the fuzzy numbers have been normalized, we already have a relative index as follows:
@@
where $\widetilde{O}_{j}^{in*}$ is the order set of fuzzy numbers with normalized support in $[0,1]$.
@@ 6.1.2 Fuzzy out-degree centrality index
When using out-degrees, it measures the gregariousness of a node.
@@
Assume that $\widetilde{G}_{dir}=\left(V,E,\widetilde{A}_{dir}\right)$ is a DFSN, $\widetilde{D}^{out}\left(v_{i}\right)$ is the sum of the fuzzy relations that are adjacent from $v_{i}$, then $\widetilde{O}_{j}^{out}\left(v_{i}\right)$ is called fuzzy out-degree centrality of $v_{i}$
For a given node $i$, let $\widetilde{O}_{j}^{out}$ be the set of fuzzy numbers $\widetilde{A}_{ij}$ ranked in decreasing order, indicating the relationship between each node $i$ and the node $j$ (directed), the formula of fuzzy out-degree centrality of $v_{i}$ is given by
@@
Naturally, a possible solution can also be to adopt a vector of decreasing weights so that the importance of fuzzy numbers decreases as the relationship is less important.
As for the in-degree centrality index, also in this case, we can define a relative out-degree centrality index as follows:
@@
6.1.3 Total Fuzzy degree centrality index
@@
However, we can consider a formula for the total degree centrality of $v_{i}$ given by:
@@ 6.2 Fuzzy betweenness centrality
Another way of looking at centrality is by considering how important nodes are in connecting other nodes. In the classical setting, for a node $v_{i}$, we can compute the relative frequency of the shortest paths to connect other nodes that pass through $v_{i}$.
In the classic case, the search for the shortest path consists of finding the path from one node to another node that is as fast as possible, crossing the least number of intermediate nodes. In our case, however, the nodes of the social network are all connected to each other with a different degree of truth; consequently, the shortest path, in a case where all the fuzzy edges would be equal, would always be the directed path from one node to another node. But usually, in reality, the connections are not all the same.
At this point in our setting, we are forced to reinterpret the concept of the shortest path as follows: In a context of fuzzy social network analysis, the importance of a node in connecting the other nodes must be evaluated based on the quality of the existing fuzzy edges: one path must be considered privileged over another if it is more effective in connecting two elements.
In our context, a path can be considered more effective than another path if it is able to connect two nodes, ensuring a higher connection strength. Clearly, it may happen that a longer path in terms of intermediate nodes to cross is better in terms of the strength of the edges and, therefore, more effective in connecting two elements $i$ to $j$. Therefore, in a network where everyone is (“more” or “less”) connected, an indirect path to go from node $i$ to $j$ could be better than the directed path from $i$ to $j$ (for example $i$ and $j$ know each other little directly but are well-connected through another subject).
@@
where $\widetilde{N}_{k}^{(1)}$ is the set of $K$ fuzzy edges in the first path between $i$ and $j$ with normalized support in $[0,1]$ and ranked in decreasing order, $\widetilde{N}_{t}^{(2)}$ is the set of fuzzy numbers in the second path between $i$ and $j$ also with normalized support in $[0,1]$ and ranked in decreasing order, and $w_{k}$ and $w_{t}$ is a vector of weights. We define $P_{i\rightarrow j}^{(1)}$ as the best path from $i$ to $j$ when for any other path $P_{i\rightarrow j}^{(2)}$ from $i$ to $j$ we have that $P_{i\rightarrow j}^{(1)}\succ P_{i\rightarrow j}^{(2)}$. Equation 31 allows us to consider different possibilities for selecting the so-called best path, e.g. average fuzzy number, minimum, maximum, etc. From now on, we will use the minimum by introducing $w_{k}=\{0,0,...,1\}$ and $w_{t}=\{0,0,...,1\}$ as the vectors of weights. In other words, when looking at two possible alternative paths to reach $j$ starting from the node $i$, we examine all the fuzzy edges in those paths, and we select the best one with the higher minimum fuzzy number (edge) in the path.
An example of this reasoning is illustrated in Figure 3. The best path for going from Dr. Perry to Dr. Porreca is via Prof. Smith. In fact, if we consider the direct connection, we have a very weak relationship, while if we connect the two nodes indirectly, we have two very strong relationships. In other words, if the direct relationship is weak, it is better to take more steps characterized by strong relationships. In the case of the example shown, the same applies if Dr. Porreca wants to get in touch with Perry (reverse path). This graphic example helps to better understand that in our context where the links are blurred, the concept of the shortest path is replaced by the concept of the best and most effective path to connect two distinct nodes. So we can say that the concept of betweenness centrality we have in the classic social network, that is, that an actor is highly central when he is in the middle of many shorter paths to go from a generic $j$ node to a generic $i$ node, in this context, it is replaced by the following concept: a node $z$ has high centrality if, considering all the best paths to go from a generic node $j$ to a generic node $i$, $z$ is frequently in the middle of these paths.
Figure 3: Example of fuzzy social network where the directed path between two nodes is worse than an indirected path in terms of fuzzy numbers edges. In other words, if the direct relationship is weak it is better to take more steps characterized by strong relationships.
Having identified the concept of the best path, we can therefore formalize the betweenness centrality of node $i$ as the ratio between all the best paths to join two nodes $s$ and $t$ via $i$ (obviously they are indirect paths), and all the best paths between all the pairs of nodes $s$ and $t$ (direct or indirect paths via $i$ or not). Formally, we can write as follows:
@@
where $P_{st}(v_{i})$ indicate the number of “best” paths between $s$ and $t$ that passes thought $i$, and $P_{st}$ the total number of “best” paths between generic $s$ and $t$.
The search of the so-called “best” paths is done according to the minimum fuzzy number edge occurring in the path between $s$ and $t$, even if it is indirect. Clearly, to compare the multiple different paths that occur in a social network, we have to rank them according to the mentioned criteria to order fuzzy numbers, e.g. the center of gravity approach.
@@ 6.3 Fuzzy closeness centrality
In classical closeness centrality, the intuition is that the more central nodes are, the more quickly they can reach other nodes. Formally, these nodes should have a smaller average shortest path length to other nodes. The smaller the average shortest path length, the higher the centrality for the node.
As we have seen for fuzzy betweenness centrality, also in this case we refer to the concept of “shortest” path as “best” path, i.e. to reach $j$ starting from $i$, one path could better than another one despite the number of intermediate nodes is higher when the relationship between them is strong.

### v2

@@ 3.4 Fuzzy Graphs
Assume $G_{dn}=(V,E_{dn})$ is a directed fuzzy social network (DFSN). The directed fuzzy connected intensity matrix $\widetilde{C}_{d}$ is a matrix where each element $\widetilde{s}(v_{i},v_{j})$ represents the directed fuzzy connected intensity from node $v_{i}$ to node $v_{j}$:
$\widetilde{C}_{d}=\begin{pmatrix}\widetilde{s}(v_{1},v_{1})&\cdots&\widetilde{s}(v_{1},v_{n})\\ \vdots&\ddots&\vdots\\ \widetilde{s}(v_{n},v_{1})&\cdots&\widetilde{s}(v_{n},v_{n})\end{pmatrix}$
The directed fuzzy connected intensity matrix is vital for determining the relationships (direct or indirect) between any two actors in the DFSN. Hereafter, we will focus on TFNs using a the following compact notation: $\tilde{A}=(a_{l},a_{m},a_{r}).$
4 Novel Fuzzy Centrality Indexes
Centrality indices assess a node’s importance in a network, reflecting the role of the corresponding entity in the modeled system. In the following sections, we extend the classical centrality indexes to the case of a network where the edges are represented by triangular fuzzy numbers, i.e. ties are vague; thus, we define fuzzy in-degree centrality, fuzzy out-degree centrality, fuzzy total-degree centrality, fuzzy betweenness centrality, fuzzy in-closeness centrality, fuzzy out-closeness centrality, and fuzzy total-closeness centrality.
@@ 4.1 Fuzzy degree centrality
In real-world interactions, we often consider people with many connections to be important. Degree centrality transfers the same idea into a measure. The degree centrality measure ranks nodes with more connections higher in terms of centrality. In the classical framework, the degree centrality $C_{d}(v_{i})$, in directed graphs can be computed in different ways according to the number of in-connections and out-connections, leading to the in-degree index $d_{i}^{in}$ and the out-degree index $d_{i}^{out}$, respectively, or the combination of both [^1] The number of edges connected to a vertex is referred to as its “degree” in graph theory. The degree of a node $v_{i}$ is commonly denoted as $d_{i}$. In the context of directed graphs, nodes have both in-degrees (edges pointing towards the node) and out-degrees (edges pointing away from the node), denoted as $d_{i}^{in}$ and $d_{i}^{out}$, respectively. For example, in social platforms like Facebook, the degree of a user corresponds to the number of friends [Zafarani et al., 2014].
@@
[…]
However, in the context of (directed) fuzzy social networks, the edges are represented by fuzzy numbers, and thus they are not scalars.
@@ 4.1.1 Fuzzy in-degree centrality index
$\widetilde{D}_{in}\left(v_{i}\right)=\bigoplus_{j\in[n],j\neq i}\left(w_{j}\odot\tilde{O}_{j}^{in}\right)$ (16)
where Equation 16 makes use of the FTOWA defined in Equation 8. To rank fuzzy numbers, we use the centre of gravity approach [^2] The Center of Gravity (CoG) of a fuzzy number $\mu(x)$ is calculated as follows: $\operatorname{CoG}(\mu)=\frac{\int_{-\infty}^{\infty}x\cdot\mu(x)\,\mathrm{d}x}{\int_{-\infty}^{\infty}\mu(x)\,\mathrm{d}x}.$ (17) where $\mu(x)$ is the membership function of the FN $\mu$, defined on the real numbers $\mathbb{R}$ with values in the interval $[0,1]$; $x$ represents the variable over which the integration is performed, encompassing all possible values within the support of the FN; the numerator $\int_{-\infty}^{\infty}x\cdot\mu(x)\,\mathrm{d}x$ computes the first moment (weighted average) of the FN, considering the distribution of $\mu(x)$ over $\mathbb{R}$, and finally, the denominator $\int_{-\infty}^{\infty}\mu(x)\,\mathrm{d}x$ represents the total area under the membership function, which normalises the moment to yield the CoG. The CoG is particularly useful in ranking FNs because it provides a real-valued metric that reflects the balance point of the FN. This method is especially effective when comparing TFNs, which are defined by their left endpoint $a_{l}$, core $a_{m}$, and right endpoint $a_{r}$. The CoG value offers an intuitive means of comparing the magnitude of different FNs within the same context. While other ranking methods exist, such as those based on medians or alternative moments, the CoG is favored in this study for its simplicity and alignment with the properties of TFNs. . It follows that we can obtain the following extreme cases:
@@
Equation 16 does not allow us to compare the centrality values between different networks. We can normalise the index to become a pure number between 0 and 1 to overcome this drawback. The reasoning behind the normalisation is that we calculate the maximum that the in-degree centrality index can take. The maximum of this index would be obtained if an actor is perfectly connected with all the other nodes. For this aim, the edge represented by the fuzzy number should have the characteristic of possessing a core with the maximum possible value on the support of the fuzzy number (with zero fuzziness). If the latter happens for all the actor’s edges, we can compute the maximum of Equation 16 as follows.
The easiest way to find this relative index is to normalise the fuzzy numbers. In fact, unlike the classical case in which the number of edges is necessary to normalise the in-degree centrality, in this case, all the nodes are connected, and therefore, the intensity of the edges matters. Once the fuzzy numbers have been normalised, we already have a relative index as follows:
@@
where $\widetilde{O}_{j}^{in*}$ is the order set of fuzzy numbers with normalised support in $[0,1]$.
@@ 4.1.2 Fuzzy out-degree centrality index
Out-degrees quantify the extent of a node’s outgoing connections, measuring its influence or activity within the network. This metric reflects how actively a node interacts with others, indicating its level of engagement or gregariousness within the network’s structure.
@@
Assume that $\widetilde{G}_{dir}=\left(V,E,\widetilde{A}_{dir}\right)$ is a DFSN, $\widetilde{D}^{out}\left(v_{i}\right)$ is the sum of the fuzzy relations that are adjacent from $v_{i}$, then $\widetilde{O}_{j}^{out}\left(v_{i}\right)$ is called fuzzy out-degree centrality of $v_{i}$.
For a given node $i$, let $\widetilde{O}_{j}^{out}$ be the set of fuzzy numbers $\widetilde{A}_{ij}$ ranked in decreasing order, indicating the relationship between each node $i$ and the node $j$ (directed), the formula of the fuzzy out-degree centrality of $v_{i}$ is given by:
@@
A possible solution is to adopt a vector of decreasing weights so that the importance of fuzzy numbers decreases as the relationship is less important. As for the in-degree centrality index, also in this case, we can define a relative out-degree centrality index as follows:
@@
4.1.3 Fuzzy total-degree centrality index
@@
However, we can consider a formula for the fuzzy total-degree centrality of $v_{i}$ given by:
@@ 4.2 Fuzzy betweenness centrality
Another way of looking at centrality is by considering how important nodes are in connecting other nodes. In the classical setting, for a node $v_{i}$, we can compute the relative frequency of the shortest paths333 A walk from node $v_{0}$ to $v_{k}$ in a graph $G=(V,E)$ is an alternating sequence $v_{0},e_{1},v_{1},e_{2},\ldots,e_{k},v_{k}$ of vertices and edges, where $e_{i}$ is $\{v_{i-1},v_{i}\}$ in the undirected case and $(v_{i-1},v_{i})$ in the directed case, with braces indicating an undirected edge and parentheses representing an ordered pair for a directed edge. The length of the walk is defined as the number of edges on the walk. The walk is called a path if $e_{i}\neq e_{j}$ for $i\neq j$, and a path is a simple path if $v_{i}\neq v_{j}$ for $i\neq j$. For a path $p$ in a graph $G=(V,E)$ with edge weights $w$, the weight of the path, denoted by $w(p)$, is defined as the sum of the weights of the edges on $p$. A path from $u$ to $v$ in $G$ is the shortest path (with respect to $w$) if its weight is the smallest possible among all paths from $v_{i}$ to $v_{j}$. to connect other nodes that pass through $v_{i}$. The search for the shortest path consists of finding the path from one node to another that is as fast as possible, crossing the least number of intermediate nodes [^4] For a node $v_{i}$, one approach is to quantify the number of shortest paths between other nodes that pass through $v_{i}$: $C_{b}(v_{i})=\sum_{\begin{subarray}{c}s\neq v_{i}\neq t\end{subarray}}\frac{\sigma_{st}(v_{i})}{\sigma_{st}}$ (22) where $\sigma_{st}$ denotes the total number of shortest paths from node $s$ to $t$ (also referred to as information pathways), and $\sigma_{st}(v_{i})$ represents the number of those paths that pass through $v_{i}$. Essentially, this metric evaluates how central $v_{i}$ is in facilitating connections between any pair of nodes $s$ and $t$. To standardize betweenness centrality across networks, normalization is necessary as follows: $C_{b}^{*}(v_{i})=\frac{C_{b}(v_{i})}{\max_{s\neq t}C_{b}(v_{i})}$ (23) where the maximum value occurs when $v_{i}$ lies on all shortest paths from $s$ to $t$ for every pair $(s,t)$, i.e., $\forall(s,t),s\neq t\neq v_{i},\frac{\sigma_{st}(v_{i})}{\sigma_{st}}=1$. .
In fuzzy social networks, where fuzzy numbers represent connections between nodes, each node is connected to every other node with varying degrees of truth. As a result, if all fuzzy connections were equal, the shortest path between two nodes would be the direct connection between them. However, in real-world scenarios, the connections typically vary in strength, leading to different shortest paths depending on the specific degrees of truth associated with each connection. Therefore, in our setting, we must reinterpret the concept of the shortest path as follows: the importance of a node in connecting other nodes must be evaluated based on the quality of the existing fuzzy edges, and thus, one path is privileged over another if more effective in connecting two elements of the network. In our context, a path can be considered more effective than another path if it can connect two nodes, ensuring a higher connection strength (it may happen that a longer path, in terms of intermediate nodes to cross, is better in terms of the strength of the edges and, therefore, more effective in connecting two elements $i$ and $j$). Hence, in a network where everyone is “more” or “less” connected, an indirect path to go from node $i$ to $j$ could be better than the directed path from $i$ to $j$ (for example, $i$ and $j$ may have a weak direct connection but are strongly connected through a mutual acquaintance).
@@
where $\widetilde{N}_{k}^{(1)}$ is the set of $K$ fuzzy edges in the first path between $i$ and $j$ with normalized support in $[0,1]$ and ranked in decreasing order, $\widetilde{N}_{t}^{(2)}$ is the set of fuzzy numbers in the second path between $i$ and $j$ also with normalized support in $[0,1]$ and ranked in decreasing order, and $w_{k}$ and $w_{t}$ are vectors of weights. We define $P_{i\rightarrow j}^{(1)}$ as the best path from $i$ to $j$ when for any other path $P_{i\rightarrow j}^{(2)}$ from $i$ to $j$ we have that $P_{i\rightarrow j}^{(1)}\succ P_{i\rightarrow j}^{(2)}$. Equation 24 allows us to consider different possibilities for selecting the so-called best path, e.g. average fuzzy number, minimum, maximum, etc. From now on, we will use the minimum by introducing $w_{k}=\{0,0,...,1\}$ and $w_{t}=\{0,0,...,1\}$ as the vectors of weights. In other words, when looking at two possible alternative paths to reach $j$ starting from the node $i$, we examine all the fuzzy edges in those paths, and we select the best one with the higher minimum fuzzy number (edge) in the path.
An example that can help clarify the reasoning behind the proposed idea is illustrated in Figure 2. The best path from Dr. Perry to Dr. Porreca is via Prof. Smith. If we consider the direct connection, we have a very weak relationship, while if we connect the two nodes indirectly, we have two very strong relationships. In other words, if the direct relationship is weak, it is better to take more steps characterised by strong relationships. In the case of the example shown, the same applies if Dr. Porreca wants to get in touch with Perry (reverse path). This graphic example helps to better understand that in our context, where the links are blurred, the concept of the shortest path is replaced by the idea of the "best" and most effective path to connect two distinct nodes.
Figure 2: Example of fuzzy social network where the directed path between two nodes is worse than an indirect path in terms of fuzzy numbers edges. In this context, the traditional concept of betweenness centrality in classic social networks, where an actor is considered highly central if they frequently lie on the shortest paths between two nodes, is redefined. Here, a node $z$ is considered highly central if it frequently appears on the optimal paths between any generic nodes $j$ and $i$.
Given the concept of the best path, we can define the fuzzy betweenness centrality as follows:
@@
where $P_{st}(v_{i})$ indicate the number of “best” paths between $s$ and $t$ that passes thought $i$ (they are indirect paths), and $P_{st}$ is the total number of “best” paths between $s$ and $t$, i.e. all the best paths between all the pairs of nodes $s$ and $t$ (direct or indirect paths via $i$ or not via $i$).
The search of the so-called “best” paths is done according to the minimum fuzzy number (edge) occurring in the path between $s$ and $t$, even if it is indirect. To compare the multiple different paths occurring in a social network, we rank them using the CoG (see Equation 17).
@@ 4.3 Fuzzy closeness centrality
In classical closeness centrality, the intuition is that the more central nodes, the more quickly they can reach other nodes. Formally, these nodes should have a smaller average shortest path length to other nodes. The smaller the average shortest path length, the higher the centrality for the node  [^5] Closeness centrality measures the average distance from a vertex to all other vertices in the network. Unlike other network metrics, closeness centrality offers a distinct perspective on each individual’s network position, capturing how quickly each vertex can reach all others [Hansen et al., 2020]. The central idea is that nodes with higher closeness centrality are more central, allowing them to reach other nodes more efficiently [Zafarani et al., 2014]. Mathematically, closeness centrality $C_{c}(v_{i})$ for vertex $v_{i}$ is defined as: $C_{c}(v_{i})=\frac{1}{\bar{l}_{D_{j}}}$ (26) where $\bar{l}_{D_{j}}=\frac{1}{n-1}\sum_{v_{j}\neq v_{i}}l_{i,j}$ is the average shortest path length from node $v_{i}$ to all other nodes in the network. Nodes with smaller average shortest path lengths $l_{i,j}$ exhibit higher closeness centrality.
As we have observed with fuzzy betweenness centrality, the concept of the "shortest" path is replaced by the "best" path. This means that, when traveling from $i$ to $j$, one path might be preferable to another, even if it involves more intermediate nodes.
This reinterpretation naturally extends to fuzzy closeness centrality, where both in-closeness and out-closeness centrality are taken into account. Fuzzy in-closeness centrality reflects how central a node is based on its accessibility from other nodes in the network. A node with high fuzzy in-closeness centrality is characterized by strong incoming connections, indicating that it can be easily reached from various parts of the network. On the other hand, fuzzy out-closeness centrality assesses a node’s centrality based on its ability to connect to other nodes. A node with high fuzzy out-closeness centrality efficiently reaches many other nodes, underscoring its influence and reach within the network. Thus, fuzzy closeness centrality provides a more nuanced understanding of a node’s position, considering not just the shortest paths but the best paths. This approach allows for a richer analysis of centrality in networks where connections differ in strength and significance.